# Unlocking the Potential of ChatGPT: A Comprehensive Exploration of its Applications, Advantages, Limitations, and Future Directions in Natural Language Processing

### v10

@@ Abstract
Large language models have revolutionized the field of artificial intelligence and have been used in various applications. Among these models, ChatGPT (Chat Generative Pre-trained Transformer) has been developed by OpenAI, it stands out as a powerful tool that has been widely adopted. ChatGPT has been successfully applied in numerous areas, including chatbots, content generation, language translation, personalized recommendations, and even medical diagnosis and treatment. Its success in these applications can be attributed to its ability to generate human-like responses, understand natural language, and adapt to different contexts. Its versatility and accuracy make it a powerful tool for natural language processing (NLP). However, there are also limitations to ChatGPT, such as its tendency to produce biased responses and its potential to perpetuate harmful language patterns. This article provides a comprehensive overview of ChatGPT, its applications, advantages, and limitations. Additionally, the paper emphasizes the importance of ethical considerations when using this robust tool in real-world scenarios. Finally, This paper contributes to ongoing discussions surrounding artificial intelligence and its impact on vision and NLP domains by providing insights into prompt engineering techniques.
@@ 2 Overview of ChatGPT and its capabilities
ChatGPT (Generative Pre-trained Transformer) is generative AI model (GAI) based on natural language processing techniques, developed by OpenAI, a leading AI research organization founded by a group of technology luminaries including Elon Musk, Sam Altman, Greg Brockman, and Ilya Sutskever [5]. This AI tool has achieved 100 million users in just three months.
@@
Over the following years, OpenAI continued to develop and refine ChatGPT, releasing several larger and more advanced versions of the model. In May 2020, OpenAI released GPT-3, the latest and most powerful version of ChatGPT, which has been widely hailed as a breakthrough in NLP. GPT-3 has over 175 billion parameters, making it the largest NLP model to date, and has been used in a range of applications, from chatbots and virtual assistants to content generation and language translation. On March 14th, 2023, OpenAI announced the release of GPT-4 [7, 8], which has 100 trillion parameters that offer greater improvements compared to those found in GPT-3 such as accepting image and text inputs and emitting text outputs. GPT-4 performs better than previous large language models and most current advanced systems on traditional NLP benchmarks. In addition, GPT-4 shows remarkable results on the MMLU benchmark, which includes multiple-choice questions covering 57 topics in English and other languages. GPT-4 not only surpasses existing models in English by a significant margin but also exhibits strong performance in other languages. On the translated versions of the MMLU benchmark, GPT-4 outperforms the state-of-the-art models presented in [9]. Recently, it has been proven that GPT-4 becomes more accurate when it adopts self-supervising learning with 30% of performance boost using the "Reflexion" technique. This means that the existing impressive capability of GPT-4 to carry out different tests is enhanced by a newly introduced framework that enables AI agents to imitate human-like self-reflection and self-evaluation [10, 11]. Essentially, it involves additional steps in which GPT-4 creates tests to analyze its own answers, identifying mistakes and weaknesses, and subsequently revises its solutions accordingly. As a result, this framework enhances the ability of GPT-4 to perform various tests more effectively as shown in Figure 3.
@@ 3 Transformers and pre-trained language models
The Transformer architecture is used as the main structure for several cutting-edge models, including GPT-3 [12], DALL-E-2 [13] and Codex [14]. It was created to overcome the limitations of traditional models like RNNs in managing sequences of varying lengths and contextual information. The Transformer relies on a self-attention mechanism, which enables the model to focus on different segments of the input sequence. The Transformer comprises an encoder and a decoder. The encoder processes the input sequence and produces hidden representations, while the decoder uses these hidden representations to generate the output sequence. Each layer of the encoder and decoder contains a multi-head attention mechanism and a feed-forward neural network. The multi-head attention is the most important part of the Transformer, as it determines how tokens are weighted based on their relevance. This method of information routing allows the model to better handle long-term dependencies, leading to improved performance across various NLP tasks. Another advantage of the Transformer is its parallelizability, which allows it to handle large-scale pre-training and adaptability to different downstream tasks without inductive biases. Google recently unveiled Gemma, a family of lightweight open models designed as a competitor to ChatGPT. Gemma serves as the foundation for the Gemini models, which are based on a transformer decoder architecture. It offers two sizes of models with either 2 billion or 7 billion parameters [15]. The release includes both pretrained and fine-tuned checkpoints for users. Multi-Query Attention [16], RoPE Embeddings [17], GeGLU Activations [18], and RMSNorm [19] are some improvements used to build the Gemma model. In April 2024, RecurrentGemma [20] that uses the novel Griffin architecture of Google is introduced with fewer trained tokens than Gemma-B2. Figure 10 presents the architecture of the transformer which is the main component of all the aforementioned models.
@@
Masked language modeling: used in models like BERT [21] and RoBERTa [22], involves predicting the probability of a masked token given its context within a sentence.
Autoregressive language modeling: used in models like GPT-3 and Open pre-trained transformer language models (OPT) [23], involves modeling the probability of the next token in a sentence given the preceding tokens, making it a left-to-right language modeling approach. Autoregressive models are better suited for generative tasks than masked language models. RoBERTa uses the same architecture as BERT but performs better by increasing the amount of pre-training data and incorporating more challenging pre-training objectives. XL-Net [24], another model based on BERT, uses permutation operations to change the prediction order for each training iteration, allowing the model to learn more information across tokens. Figure 5 depicts the timeline of the most known generative models with their number of parameters.
@@ 4.1 Healthcare:
  • Telemedicine: AI can support remote patient consultations and improve healthcare accessibility in remote areas.
@@
ChatGPT can be used to improve the quality of the previous services by providing better and more accurate help to the assistants. For example, Figure 7 shows the answer of ChatGPT about Covid-19 symptoms. These assistants can provide personalized recommendations and advice, such as medication reminders or diet plans, and help patients manage their health conditions. Moreover, ChatGPT has the potential to generate discharge summaries, which are comprehensive medical documents summarizing a patient’s hospital stay. More Potential Benefits of ChatGPT in Healthcare can be found in [26]. Wang et al. [27] proposed a study to compare the level of knowledge and interpretation skills between ChatGPT and medical students in China. To achieve this goal, the Chinese National Medical Licensing Examination was administered to both ChatGPT and medical students, and their performances were compared. Johnson et al. [28] assessed the accuracy and reliability of ChatGPT to generate medical responses. In this study, a total of 33 physicians from 17 different specialties were involved in the generation of 284 medical questions. These questions were then categorized by the physicians themselves based on their perceived difficulty level, which was classified as either easy, medium, or hard. The questions were designed to elicit either binary (yes/no) responses or descriptive answers. On the other hand, Benoit [29] has tested ChatGPT to assess its potential in quickly generating, rewriting, and evaluating sets of clinical vignettes. The evaluation included diagnostic and triage accuracy of many diseases including anaphylaxis, asthma, bronchiolitis, croup, ear infection, fever, functional constipation, and gastroenteritis. Similarly, Gunther Eysenbach[30] conducted experiments using ChatGPT to explore possible applications of chatbots in medical education. ChatGPT demonstrated its ability to create a virtual patient simulation and quizzes for medical students, evaluate a simulated doctor-patient interaction, and even summarize a research article. Usually, junior doctors are responsible for composing these summaries, but due to their workload, they may not be given the priority they deserve, causing delays in patients’ discharges and incomplete summaries [31]. This puts pressure on an already overworked junior doctor workforce and may lead to potential patient safety issues during the transition from secondary to primary care. However, if ChatGPT is utilized, it could alleviate the burden of composing discharge summaries, resulting in more timely and high-quality summaries that are essential for safe patient care [32]. Other ChatGPT decision-making and clinical decision support can be found in [33, 34, 35, 36, 37, 38, 39, 40].

### v11

@@ Abstract
Large language models, pivotal in artificial intelligence, find diverse applications. ChatGPT (Chat Generative Pre-trained Transformer), an OpenAI creation, stands out as a widely adopted, powerful tool. It excels in chatbots, content generation, language translation, recommendations, and medical applications, due to its ability to generate human-like responses, comprehend natural language, and adapt contextually. Its versatility and accuracy make it a potent force in natural language processing (NLP). Despite successes, ChatGPT has limitations, including biased responses and potential reinforcement of harmful language patterns. This article offers a comprehensive overview of ChatGPT, detailing its applications, advantages, and limitations. It also describes the main advancements from GPT-3 to GPT-4 Omni, comparing them with other LLMs like LLaMA 3 and Gemini. The paper underscores the ethical imperative when utilizing this robust tool in practical settings. Furthermore, it contributes to ongoing discussions on artificial intelligence’s impact on vision and NLP domains, providing insights into prompt engineering techniques.
@@ 2 Overview of ChatGPT and its capabilities
ChatGPT (Generative Pre-trained Transformer) is generative AI model (GAI) based on natural language processing techniques, developed by OpenAI, a leading AI research organization founded by a group of technology luminaries including Elon Musk, Sam Altman, Greg Brockman, and Ilya Sutskever [5]. This AI tool has achieved 100 million users in just three months. We differentiate between two main evolutions of ChatGPT based on the functionalities it offers, which are either focused solely on NLP or incorporate visual information. We also compare the recent advancement of ChatGPT with LLaMa 3 and Gemini.
NLP-based:
@@
Over the following years, OpenAI continued to develop and refine ChatGPT, releasing several larger and more advanced versions of the model. In May 2020, OpenAI released GPT-3, the latest and most powerful version of ChatGPT, which has been widely hailed as a breakthrough in NLP. GPT-3 has over 175 billion parameters, making it the largest NLP model to date, and has been used in a range of applications, from chatbots and virtual assistants to content generation and language translation. On March 14th, 2023, OpenAI announced the release of GPT-4 [7, 8], which has 100 trillion parameters that offer greater improvements compared to those found in GPT-3 such as accepting image and text inputs and emitting text outputs.
@@
GPT-4 performs better than previous large language models and most current advanced systems on traditional NLP benchmarks. In addition, GPT-4 shows remarkable results on the MMLU benchmark, which includes multiple-choice questions covering 57 topics in English and other languages. GPT-4 not only surpasses existing models in English by a significant margin but also exhibits strong performance in other languages. On the translated versions of the MMLU benchmark, GPT-4 outperforms the state-of-the-art models presented in [9]. Recently, it has been proven that GPT-4 becomes more accurate when it adopts self-supervising learning with 30% of performance boost using the "Reflexion" technique. This means that the existing impressive capability of GPT-4 to carry out different tests is enhanced by a newly introduced framework that enables AI agents to imitate human-like self-reflection and self-evaluation [10, 11]. Essentially, it involves additional steps in which GPT-4 creates tests to analyze its own answers, identifying mistakes and weaknesses, and subsequently revises its solutions accordingly. As a result, this framework enhances the ability of GPT-4 to perform various tests more effectively as shown in Figure 3.
NLP and visual-based:
GPT-4o ("o" stands for "omni") introduced by Open AI on May 14, 2024, represents a significant advancement in creating more natural human-computer interactions. It can take input in various forms—text, audio, images, and video—and generate outputs in text, audio, or images [12]. Impressively, it can respond to audio inputs in as little as 232 milliseconds, with an average response time of 320 milliseconds, which is comparable to human response times in conversation. In terms of performance, GPT-4o matches GPT-4 Turbo for text processing in English language teaching [13] and coding tasks, while offering enhanced capabilities in non-English text processing. Additionally, it is notably faster and 50% more cost-effective in the API. GPT-4o also excels in understanding vision and audio compared to previous models. Below, we highlight the key enhancements of GPT-4o compared to previous versions.
  1. 1.
    Multimodal Integration: GPT-4o can process and generate text, images, and audio, allowing for richer and more interactive responses. This makes it versatile for various applications, including education, customer support, and content creation [14, 15].
  2. 2.
    Improved Contextual Awareness: It maintains context over longer conversations, providing more coherent and contextually relevant responses. This enhancement is particularly beneficial for personal assistants and customer support systems [14].
  3. 3.
    Enhanced Language Understanding: Trained on a larger and more diverse dataset, GPT-4o offers improved accuracy and coherence in text generation. This makes it more effective in handling complex conversations and nuanced topics [14].
  4. 4.
    Advanced Tokenization: GPT-4o features a better tokenization model, especially for non-Roman alphabets, which reduces the number of tokens needed and speeds up text generation. This also makes the model more cost-efficient [15].
  5. 5.
    Cost-Efficiency: The model is optimized to be 50% more cost-efficient than its predecessors, making sophisticated AI tools more accessible to a broader range of users [16].
  6. 6.
    Real-Time Collaboration: GPT-4o supports real-time collaboration features, allowing multiple users to interact with the AI simultaneously, which is useful in team environments and project management [14].
  7. 7.
    Customization and Fine-Tuning: Users can tailor the model to better suit specific applications by adjusting its behavior, tone, and response style, ensuring that it aligns more closely with unique requirements [14].
  8. 8.
    Desktop Application: Alongside the launch of GPT-4o, OpenAI introduced a ChatGPT desktop app, enhancing usability and integration into various workflows [15].
From the above enhancements compared to previous versions, GPT-4o can impressively analyze and interpret visual information from images and videos, enabling it to provide detailed descriptions, recognize objects, and even understand complex scenes. This multimodal capability enhances its usability in various applications such as security, content moderation, and educational tools. Visual capabilities are described in the following.
  • Image Analysis: When provided with a photo, GPT-4o can describe the contents of the image, identify objects, and even detect emotions from facial expressions. This makes it useful for applications like social media content moderation, automated tagging, and accessibility tools for visually impaired users.
  • Video Analysis: GPT-4o can process video inputs from cameras, offering functionalities like real-time object tracking, action recognition, and scene understanding. This enhancement is valuable in security systems, where the AI can monitor video feeds for unusual activities or potential threats, and in automated video editing tools that can identify key moments in footage.
  • Educational Applications: In educational settings, GPT-4o’s ability to incorporate visual aids into its responses can significantly enhance learning. For instance, when explaining a scientific concept, it can include relevant diagrams or video snippets, making complex ideas easier to grasp.
  • Healthcare: In healthcare, GPT-4o’s visual analysis capabilities can assist in diagnosing medical images, such as X-rays or MRIs, by identifying anomalies and providing preliminary assessments, thereby aiding medical professionals.
  • Custom Visual Model Training: Users can fine-tune GPT-4o’s visual processing abilities to suit specific needs. For instance, a business could train the model to recognize their products in images or videos, streamlining inventory management and quality control processes.
The introduction of GPT-4o marks a pivotal moment in AI development, combining increased accessibility, improved performance, and enhanced multimodal capabilities to push the boundaries of human-machine interaction [14, 16].
Comparison with LLaMA 3:
LLaMA 3, the third iteration of the LLaMA (Large Language Model Meta AI) series, was introduced by Meta (formerly Facebook) in 2024. It builds on the advancements of its predecessors, LLaMA 1 and LLaMA 2, incorporating improvements in model architecture, training techniques, and scalability. LLaMA 3 is designed to enhance performance across a variety of natural language processing tasks and to address some of the limitations observed in earlier models with 70B parameter scale. While GPT-3 improves upon GPT-2 primarily by increasing the amount of training data, LLaMA-3 follows a more extensive development trajectory. Unlike the incremental improvements seen from GPT to GPT-2 to GPT-3, LLaMA-3 represents a broader evolution in model design, incorporating more complex and capable configurations [17, 18]. The development of LLaMA-3 involves multiple stages, including development, training, testing, and fine-tuning across various tasks. These tasks utilize a combination of supervised learning with and without reward functions, transfer learning, and reinforcement learning (RL) techniques [19]. The model also introduces incremental modifications to the GPT architecture, enhancing its scalability and performance.
OpenAI’s original GPT, released in 2016, was designed for supervised and semi-supervised training using large datasets, with fine-tuning achievable with minimal examples. It employed a transformer architecture based on attention mechanisms, which proved effective for processing human language and complex tasks. In contrast, LLaMA-3 utilizes a different transformer-based model known as the Linear Modular Unit (LMU), which was initially designed for control systems but later adapted for human language processing [20, 21, 22]. The LMU combines the strengths of successful language models while offering scale-invariance similar to Long Short-Term Memory (LSTM) networks, making it highly efficient for handling increasingly complex language tasks. Table 1 provides a comparative analysis of several LLMs across various benchmarks and shot configurations. It includes five models: GPT-3.5, GPT-4, PaLM, PaLM-2-L, and LLaMA 2. This table highlights that GPT-4 generally performs better across most benchmarks, while models like PaLM-2-L and LLaMA 2 show strong performances in specific tasks.
Table 1: Benchmark comparison for various LLM models.
Benchmark (shots) GPT-3.5 GPT-4 PaLM PaLM-2-L LLaMA 2
MMLU (5-shot) 70.0 86.4 69.3 78.3 68.9
TriviaQA (1-shot) - - 81.4 86.1 85.0
Natural Questions (1-shot) - - 29.3 37.5 33.0
GSM8K (8-shot) 57.1 92.0 56.5 80.7 56.8
HumanEval (0-shot) 48.1 67.0 26.2 - 29.9
BIG-Bench Hard (3-shot) - - 52.3 65.7 51.2
Comparison with Gemini:
The Gemini model, developed by Google DeepMind, represents a significant advancement in the field of large language models. Four distinct versions of Gemini have been identified: Ultra, Pro, Flash, and Nano introduced on December 6, 2023, each tailored to different performance and application needs [23]. Originally known as "Gemini" during its development phase, it was officially released as Gemini 1, reflecting Google’s ongoing innovation in artificial intelligence. The Gemini model is part of Google’s strategic efforts to maintain a leading position in AI development, competing directly with other top LLMs such as OpenAI’s GPT series and Meta’s LLaMA models.
Gemini is built on a sophisticated architecture designed to enhance its ability to understand and generate human-like text across a wide range of natural language processing tasks, including text generation, summarization, machine translation, and conversational AI [24]. The model emphasizes advanced reasoning capabilities, allowing it to tackle complex tasks that require a deep understanding of context and nuance. One of Gemini’s key features is its focus on improving response accuracy and reducing biases, a common challenge in earlier LLMs [25]. It integrates state-of-the-art techniques in AI ethics and safety, aiming to minimize harmful outputs while maintaining high performance in generating relevant and contextually appropriate responses. With its scalable and efficient architecture, Gemini 1 is optimized for processing large datasets and delivering quick, accurate responses, leveraging Google’s vast computational resources. As part of a broader initiative, Gemini is deeply embedded within Google’s ecosystem, enhancing services like Google Search, cloud platforms, and other tools where advanced natural language understanding is critical. The release of Gemini marks a pivotal moment in Google’s AI strategy, ensuring competitiveness in a rapidly advancing industry. Unlike standalone models, Gemini is expected to evolve as a foundation for future iterations, pushing the boundaries of AI’s ability to understand and interact with human language. When compared to ChatGPT and LLaMA 3, Gemini distinguishes itself with its advanced reasoning and deep integration within Google’s ecosystem. Table 2 presents a comparison between LLaMA 3, Gemini Pro and Claude 3 Sonnet models across five benchmarks: MMLU, GPQA, HumanEval, GSM-8K, and MATH. Meta Llama 3 70B generally scores higher in MMLU, HumanEval, and GSM-8K, while Gemini Pro 1.5 leads in GPQA and MATH.
While ChatGPT is widely recognized for its versatility, ease of integration, and proven reliability across diverse applications, Gemini offers a unique focus on ethical AI practices, bias reduction, and seamless functionality within Google’s expansive range of services. Both models represent the cutting edge of AI technology, yet they cater to slightly different needs within the AI landscape.
Table 2: Performance comparison of Meta Llama 3 70B, Gemini Pro 1.5, and Claude 3 Sonnet
Meta LLama 3 70B Gemini Pro 1.5 Claude 3 Sonnet
MMLU (5-shot) 82.0 81.9 79.0
GPQA (0-shot) 39.5 41.5 38.5
HumanEval (0-shot) 81.7 71.9 73.0
GSM-8K (8-shot, CoT) 93.0 91.7 92.3
MATH (4-shot, CoT) 50.4 58.5 (Minerva prompt) 40.5
@@ 3 Transformers and pre-trained language models
The Transformer architecture is used as the main structure for several cutting-edge models, including GPT-3 [26], DALL-E-2 [27] and Codex [28]. It was created to overcome the limitations of traditional models like RNNs in managing sequences of varying lengths and contextual information. The Transformer relies on a self-attention mechanism, which enables the model to focus on different segments of the input sequence (See Figure 5). The Transformer comprises an encoder and a decoder. The encoder processes the input sequence and produces hidden representations, while the decoder uses these hidden representations to generate the output sequence. Each layer of the encoder and decoder contains a multi-head attention mechanism and a feed-forward neural network. The multi-head attention is the most important part of the Transformer, as it determines how tokens are weighted based on their relevance. This method of information routing allows the model to better handle long-term dependencies, leading to improved performance across various NLP tasks. Another advantage of the Transformer is its parallelizability, which allows it to handle large-scale pre-training and adaptability to different downstream tasks without inductive biases. Google recently unveiled Gemma, a family of lightweight open models designed as a competitor to ChatGPT. Gemma serves as the foundation for the Gemini models, which are based on a transformer decoder architecture. It offers two sizes of models with either 2 billion or 7 billion parameters [29]. The release includes both pretrained and fine-tuned checkpoints for users. Multi-Query Attention [30], RoPE Embeddings [31], GeGLU Activations [32], and RMSNorm [33] are some improvements used to build the Gemma model. In April 2024, RecurrentGemma [34] that uses the novel Griffin architecture of Google is introduced with fewer trained tokens than Gemma-B2. GPT-Neo, developed by the open-source community EleutherAI, is an accessible alternative to GPT-3. Founded in 2020, EleutherAI aims to democratize AI by creating and sharing open-source models with 2.7 B parameters [35]. GPT-J [36] is an open-source large language model created by EleutherAI, designed to offer high performance similar to GPT-3. It has 6 billion parameters and is known for its ability to generate coherent and contextually relevant text. It is freely available, supporting research and development in AI.
Open-source models thrive on community collaboration, fostering rapid innovation and broader applicability through diverse contributions from developers, researchers, and enthusiasts. This collaborative environment accelerates the refinement and expansion of these models to address a wide range of problems. In contrast, closed-source models benefit from structured, corporate-backed development, which provides the advantage of faster deployment, comprehensive support, and targeted solutions tailored to specific business needs. Despite these benefits, the debate over which approach is superior persists, as each has its strengths and limitations depending on the context and objectives. Table 4 compares open-source and closed-source models across various aspects, including development model, contributors, innovation, and provides examples for each category. Figure 4 presents a comparison between the best-performing closed-source and open-source models reported by Arena Elo rating system between Mar 2023 and Apr 2024.
Figure 4: Open-source and closed-source models according to Arena Elo rating system.
Table 4: Comparison of open source and closed source models
Aspect Open source models Closed source models
Development Model Community-driven, collaborative, iterative Centralized, structured, product-focused
Contributors Developers, researchers, and enthusiasts Corporate teams with specialized roles
Innovation Rapid iteration, diverse perspectives, and unforeseen evolution Structured, focused on specific business needs
Applicability Versatile, adaptable to a wide range of problems Tailored to specific business applications and market needs
Ownership Shared among contributors, fostering investment and engagement Owned by a corporate entity with clear control over the development
Resource access Community-driven, often limited resources Backed by extensive corporate resources and technical expertise
Documentation and support Variable, depending on community contributions Comprehensive, usually includes training and support
Deployment speed Can be slower due to the need for community-driven refinement Typically faster due to structured development and corporate backing
Market readiness Requires additional effort to commercialize Generally product-validated and market-ready
Efficacy and customization High potential for efficacy due to diverse contributions and use cases High efficiency for specific business needs and integration
Customer understanding Deep understanding from contributors who may also use the models Often based on the needs of the corporation and its clients
Examples GPT-Neo, GPT-J, LLaMA, Bloom, T5 GPT-4, Bard, Claude (Anthropic), Gemini
@@
Autoregressive language modeling: used in models like GPT-3 and Open pre-trained transformer language models (OPT) [37], involves modeling the probability of the next token in a sentence given the preceding tokens, making it a left-to-right language modeling approach. Autoregressive models are better suited for generative tasks than masked language models. RoBERTa uses the same architecture as BERT but performs better by increasing the amount of pre-training data and incorporating more challenging pre-training objectives. XL-Net [38], another model based on BERT, uses permutation operations to change the prediction order for each training iteration, allowing the model to learn more information across tokens.
Masked language modeling: used in models like BERT [39] and RoBERTa [40], involves predicting the probability of a masked token given its context within a sentence. Figure 6 depicts the timeline of the most known generative models with their number of parameters.
@@ 4.1 Healthcare:
  • Telemedicine: AI supports remote patient consultations and improve healthcare accessibility in remote areas.
@@
ChatGPT can significantly enhance the quality of existing services by offering more accurate and tailored assistance to both users and service providers. By leveraging advanced natural language processing capabilities, ChatGPT can understand and respond to complex queries with precision, thereby improving the efficiency and effectiveness of customer support, virtual assistants, and other interactive services. For example, Figure 8 shows the answer of ChatGPT about Covid-19 symptoms. These assistants can provide personalized recommendations and advice, such as medication reminders or diet plans, and help patients manage their health conditions. Moreover, ChatGPT has the potential to generate discharge summaries, which are comprehensive medical documents summarizing a patient’s hospital stay. More Potential Benefits of ChatGPT in Healthcare can be found in [42]. Wang et al. [43] proposed a study to compare the level of knowledge and interpretation skills between ChatGPT and medical students in China. To achieve this goal, the Chinese National Medical Licensing Examination was administered to both ChatGPT and medical students, and their performances were compared. Johnson et al. [44] assessed the accuracy and reliability of ChatGPT to generate medical responses. In this study, a total of 33 physicians from 17 different specialties were involved in the generation of 284 medical questions. These questions were then categorized by the physicians themselves based on their perceived difficulty level, which was classified as either easy, medium, or hard. The questions were designed to elicit either binary (yes/no) responses or descriptive answers.
On the other hand, Benoit [45] has tested ChatGPT to assess its potential in quickly generating, rewriting, and evaluating sets of clinical vignettes. The evaluation included diagnostic and triage accuracy of many diseases including anaphylaxis, asthma, bronchiolitis, croup, ear infection, fever, functional constipation, and gastroenteritis. Similarly, Gunther Eysenbach[46] conducted experiments using ChatGPT to explore possible applications of chatbots in medical education. ChatGPT demonstrated its ability to create a virtual patient simulation and quizzes for medical students, evaluate a simulated doctor-patient interaction, and even summarize a research article. Usually, junior doctors are responsible for composing these summaries, but due to their workload, they may not be given the priority they deserve, causing delays in patients’ discharges and incomplete summaries [47]. This puts pressure on an already overworked junior doctor workforce and may lead to potential patient safety issues during the transition from secondary to primary care. However, if ChatGPT is utilized, it could alleviate the burden of composing discharge summaries, resulting in more timely and high-quality summaries that are essential for safe patient care [48]. Other ChatGPT decision-making and clinical decision support can be found in [49, 50, 51, 52, 53, 54, 55, 56].